OpenStudy (anonymous):

What statement makes the open sentence 49 – 7x = 21 true? A. x = 11 B. x = 7 C. x = 4 D. x = 3

OpenStudy (anonymous):

I got D. when I did it.

OpenStudy (idealist10):

49-7x=21 49-21=7x 28=7x x=28/7 x=4 Answer: C

OpenStudy (anonymous):

Thanks so much.